Find the maximum sum of 2 numbers given that the first plus the square of the second is equal to 30?
Answer by richard1234(7193)

We want to maximize given that . We can let a = 30 - b^2, so we want to maximize



Recall that the max/min of a quadratic function ax^2 + bx + c occurs at x = -b/2a. In this case, the maximum occurs at b = 1/2, which implies a = 119/4. Therefore a+b = 121/4.